Shyam is four times as old as Ram. After 7 years, Shyam's age will be thrice of Ram. shyam's present age is
1)14 years 2)48 years
3)52 years 4)56 years

Found 2 solutions by balthazar612, krish2002:
Answer by balthazar612(14)

ans d 56 years is correct

Answer by krish2002(1)

Let Rams age = x
Then Shyam age = 4x
After years,
4x+7 = 3(x+7)
4x+7 = 3x + 21
4x-3x=21-7
x = 14 (Rams Age)
Hence Shyams age = 4x14 = 56years
Ans = 56 years